### v1

# Design of Magnetic Lattices with Quantum Optimization Algorithms

Abstract
This article investigates the identification of magnetic spin distributions in ferromagnetic materials by minimizing the system’s free energy. Magnetic lattices of varying sizes are constructed, and the free energy is computed using an Ising model that accounts for spin-to-spin neighbor interactions and the influence of an external magnetic field. The problem reduces to determining the state of each spin, either up or down, leading to an optimization problem with $2^{n\times n}$ design variables for an $n\times n$ lattice. To address the high-dimensional and computationally intractable nature of this problem, particularly for large domains, we employ a quantum optimization algorithm, BQP. The BQP results are first validated against solutions obtained using a genetic algorithm for smaller lattices. Finally, the approach is extended to large-scale systems, including $50\times 50$ lattices, where conventional methods become impractical.
preprint: AIP/123-QED
I Introduction
Magnetic materials continue to captivate researchers and engineers due to their critical role in numerous technological applications, particularly in sectors such as energy and transportation. Understanding and determining the magnetic behavior on small length scales is key to maximizing functional stability and magnetic efficiency under varying operational conditions.
To model and analyze the complex behavior of magnetic systems, a variety of theoretical frameworks have been developed. These models broadly fall into two categories: those based on localized magnetic moments and those described by electronic band structures. Among the latter, the Hubbard Herrmann and Nolting (1997), Stoner Tannous and Gieraltowski (2008), and Kondo Fazekas and Müller-Hartmann (1991) models, as well as approaches rooted in Density Functional Theory (DFT) Bihlmayer (2020), provide insight into the role of electronic structure and correlations. In contrast, models such as the Heisenberg and Ising frameworks deal with discrete spin variables and their interactions, with the Ising model Ising (1925) being particularly effective in representing simplified spin systems.
In this study, the Ising model is employed to explore the landscape of spin configurations and to determine those that minimize or optimize the system’s free energy. As with other Hamiltonian-based models, the Ising framework calculates the total energy of a system as a function of spin arrangements and their interactions, serving as a foundational approach in statistical mechanics Cipra (1987). Despite its usefulness, the model’s computational demands scale rapidly with lattice size and dimensionality, posing challenges for classical simulations. These limitations have spurred significant research on numerical efficiency, including efforts to address scaling behavior, spin correlation functions, and hysteresis effects Sethna, Dahmen, and Perkovic (2004). This causes limitations on investigating the effects of the uncertainty arising from external parameters, including the temperature and external magnetic fields. It is also the reason for potentially numerically intractable solutions when sampling-based uncertainty quantification (UQ) algorithms (e.g., Monte Carlo Simulation (MCS)) are incorporated to investigate the effects of uncertainty on these systems. As solving the optimal spin configurations to determine the magnetic state of the materials under the effects of uncertainty requires a computationally demanding and expensive design-under-uncertainty solution, there is a need to shift from more traditional optimization solutions (e.g., genetic algorithm) that can address the discrete nature of this problem to more advanced computing strategies.
To overcome these challenges, optimization methods that employ quantum principles have emerged as promising alternatives. In particular, Quantum-Inspired Evolutionary Optimization (QIEO) techniques combine classical evolutionary strategies with computational structures inspired by quantum mechanics, most notably the concepts of quantum superposition and rotation gates Kumar et al. (2024). In QIEO, each design variable is encoded using a probabilistic representation called a Q-bit, which maintains a probability amplitude for both binary states Zhang (2011). The system evolves by applying quantum-inspired rotation operations that guide the search toward optimal configurations based on fitness evaluations. This probabilistic representation facilitates a balance between exploration and exploitation, improving convergence for high-dimensional discrete problems such as those found in magnetic spin lattice design Han and Kim (2002); Rylander et al. (2001); Zhang, Han, and Kim (2006).
A closely related approach is the Quantum Approximate Optimization Algorithm (QAOA) Farhi, Goldstone, and Gutmann (2014), which formulates discrete optimization problems as minimizing the expectation value of a problem-specific Hamiltonian on quantum states parameterized by a variational circuit. Although QAOA is typically run on quantum hardware, its mathematical foundation shares key principles with QIEO, notably, encoding solutions in the state vector of a quantum system and evolving them via parameterized unitaries that are analogous to update steps in QIEO. Moreover, as demonstrated by Lloyd et al. Lloyd, Garnerone, and Zanardi (2016), quantum algorithms have the potential to efficiently capture topological and geometric features in high-dimensional configuration spaces. This is particularly valuable in problems like spin lattice optimization, where the solution landscape is highly nonconvex and spatial correlations between variables (i.e., spins) play a critical role in defining system energy. These insights reinforce the applicability of BQP’s hybrid classical-quantum strategy in navigating complex magnetic design spaces with embedded physical constraints Kumar et al. (2024, 2025).
This work builds upon these insights by utilizing quantum-enhanced methods to directly identify optimal spin configurations, shifting the emphasis away from transition temperatures and toward configuration-space optimization as a central goal in magnetic material analysis.
II Ising Model Formulation
II.1 Mathematical Formulation
The Ising model describes interactions among spins, which represent the atomic magnetic moments arranged on a lattice. Beyond condensed matter physics, it has found wide applications in diverse fields, including sociology and neural networks.Sornette (2014). In this study, the Ising model is used to investigate the magnetic behavior of 2D lattices while incorporating short- and long-range order interactions between the spins.
The energy function for the Ising model is defined as Selinger (2016):
$H=-\omega_{j}\Sigma_{<i,j>}\sigma_{i}\sigma_{j}-h\Sigma_{i}\sigma_{i}$ (1)
In Eq. 1, $h$ is the external magnetic field, $\sigma$ shows the spin parameter, while the $\sigma_{i}\sigma_{j}$ term represents interactions between spins indexed by $i$ and $j$. Moreover, $\omega_{j}$ is the weight parameter defining the weighted importance of spin-to-spin interactions based on distance or order of neighboring information. Accordingly, this weight parameter correlates with the degree of neighboring information defined in the Hamiltonian energy equation.
Solving the magnetic state of the material requires the minimization of the energy function. However, when large lattices are considered, the computation of the minimum free energy requires many iterations that are potentially intractable. To address this, it is common to only consider the interactions within the nearest neighbor spins. Nevertheless, only considering the nearest-neighbor interactions has some drawbacks, as long-range interactions can play an effective role in determining the magnetic state of the material. As a result, the modeling with only nearest-neighbor interactions may overlook important contributions from long-range neighbors or external factors such as external magnetic fields. To improve the prediction accuracy for the magnetic state of the material, the long-range order interactions are accounted for by implementing a Gaussian weight formulation as discussed in the next section.
II.2 Investigation of Spin-to-Spin Interactions
Spin-to-spin correlation functions quantify how the magnetic orientation (spin) of a particle relates to that of another at a spatially distinct site within a material. These correlations provide critical insight into the extent and nature of magnetic interactions across a system. The characteristic length scale over which spins remain correlated, which is defined as the spin correlation length, serves as an effective order parameter. It is particularly useful in capturing the emergence or breakdown of long-range magnetic order Stanley (1971); Yeomans (1992). Near critical points, such as those associated with phase transitions, this length can diverge, marking significant structural or dynamical changes in the spin configuration Kardar (2007). Variations in external parameters, including temperature and applied magnetic fields, modulate these correlations and give rise to distinct magnetic phases and critical behavior. Thus, a detailed understanding of spin correlations is not only fundamental for probing the onset of phase transitions but also essential for predicting the response of magnetic systems to external perturbations.
To incorporate long-range interactions within the spin system, the interaction strength, denoted by $\omega_{j}$, is characterized through a Gaussian distribution Hasan and Acar (2022). The Gaussian formulation used to describe the spatial variation of interaction strength is presented in Eq. 2, where $d$ represents the inter-spin distance, and $\mu_{d}$ and $\sigma_{d}$ correspond to the mean and standard deviation of the distance distribution, respectively Eğer and Acar (2024); Eger and Acar (2024). Furthermore, $\mu_{d}$ and $\sigma_{d}$ are defined as temperature-dependent variables modeled by additional Gaussian functions, such that $\mu_{d}=\mu_{d}(T)$ and $\sigma_{d}=\sigma_{d}(T)$, enabling the model to capture thermally driven changes in the correlation profile across the spin lattice Eğer and Acar (2024).
$\omega=f(d,\mu_{d}(T),\sigma^{2}_{d}(T))=\frac{1}{\sqrt{2\sigma^{2}\pi}}e^{-(d-\mu)^{2}/2\sigma^{2}}$ (2)
At low temperatures, the interaction strength between spins tends to be nearly uniform, regardless of the spatial separation between them. However, as the temperature increases, this distribution shifts, which results in a diminished influence from distant neighbors relative to those in closer proximity. This attenuation of long-range interaction strength persists even beyond the phase transition, ultimately leading to a generalized weakening of spin interactions across the lattice, including those between nearest neighbors.
While the precise functional form of interaction strength across all temperature values is not explicitly defined in the literature, the behavior of the system is well-characterized by spin correlation functions under three primary regimes: above, below, and at the Curie temperature ($T_{c}$), which marks the critical point for the ferromagnetic to paramagnetic phase transition Wipf (2013). To effectively incorporate long-range interactions in our model, we construct a statistical formulation constrained to capture this behavior accurately across temperature regimes.
$\omega_{n}\leq\omega_{n-1}\leq\omega_{n-2}\leq....\leq\omega_{3}\leq\omega_{2}\leq\omega_{1}$ (3)
Here, the parameter $n$ is determined by the lattice size, while $\omega_{1}$ and $\omega_{n}$ represent the interaction strengths of the nearest and farthest neighboring spins, respectively. To account for long-range magnetic interactions, a window parameter $\omega_{n}$ is introduced. This parameter, which scales with the lattice size, ensures that all spin interactions across the lattice are captured. The window-based formulation has previously proven effective in predicting the spatio-temporal evolution of microstructures within an Ising model framework using Markov Random Fields Acar and Sundararaghavan (2019, 2016), and has also been successfully employed in Ising models to investigate the onset of the ferromagnetic–paramagnetic phase transition Eğer and Acar (2024).
II.3 Modeling the Uncertainty of Temperature and External Field
In realistic settings, temperature and external magnetic field values often exhibit uncertainty due to factors such as sensor inaccuracies, environmental disturbances, and system-level noise. To account for these sources of variability, both temperature ($T$) and magnetic field strength ($h$) are modeled as independent random variables by assuming that they agree with a Gaussian distribution with the following mean values: $T_{\text{mean}}=298\text{K}$ and $h_{\text{mean}}=5$, with each parameter subject to a standard deviation equal to 10% of its respective mean value, capturing expected deviations in experimental or operational contexts.
To efficiently sample the input space defined by these random parameters, Latin Hypercube Sampling (LHS) is employed. A total of 3,000 design samples are generated to ensure broad and stratified coverage of the input domain. Following the sampling phase, the genetic algorithm is implemented to identify the optimum spin configurations under uncertainty. The genetic algorithm is well-suited to this task due to its effectiveness in exploring high-dimensional, non-convex solution spaces, particularly in the presence of complex spin interactions and long-range correlations. The optimization problem is further elaborated in the next section.
III Optimization of Magnetic Spin States under Uncertainty
The states of magnetic spins are identified while accounting for the uncertainty of temperature and magnetic field strength. The stochastic optimization problems are defined using two different objectives. The first optimization problem aims to solve for the optimum spin distributions that minimize the expected value of the system’s free energy ($E[f_{m}]$) to detect the magnetic equilibrium state, while the second problem also considers the effects of the uncertainty of temperature and external magnetic fields by minimizing the resultant deviations in the free energy value ($std_{f_{m}}$). The problems are first solved using the genetic algorithm. By considering the computationally expensive nature of this global optimization algorithm, the problems are solved in a $10\times 10$ magnetic domain. The problems are formulated as shown in Eqs. 4 and 5, respectively:
$\min E[f_{m}]$ (4)
$\text{find }\sigma_{i,j}\ \ \ \text{where }i,j=1,2,...,10$
$\min std_{f_{m}}$ (5)
$\text{find }\sigma_{i,j}\ \ \ \text{where }i,j=1,2,...,10$
Figure 1: Solution of the first optimization problem identifying the optimum spin distributions that minimize the expected value of the system’s free magnetic energy under the effects of the uncertainty of temperature and external magnetic field using the genetic algorithm.
The optimum solution for the first problem is obtained within just a few seconds as shown in Fig. 1, so no additional optimization algorithm is required. In contrast, uncertainty significantly impacts the optimum solution for the second problem, unlike the first problem that yields the same result as the deterministic case. Because uncertainty influences the optimum design, classical algorithms such as GA require considerably more computational time. This highlights the need for a more efficient optimization algorithm.
Figure 2: Exemplary solutions of the second optimization problem identifying the optimum spin distributions that minimize the standard deviation of the magnetic free energy under the effects of the uncertainty of temperature and external magnetic field using the genetic algorithm.
The second problem is also found to lead to multiple optimum solutions, all providing the same value for the objective function, and the reported times for the genetic algorithm and BQP include the solution time for any solution shown in Fig. 2.
IV Application of Quantum Optimization Algorithm
IV.1 Quantum-Inspired Evolutionary Optimization Algorithm
BQPhy®optimization solver primarily uses the quantum-inspired evolutionary optimization algorithm. QIEO uses traditional HPCs and is developed on the basis of the principles of quantum computing. It primarily uses concepts such as qubits, quantum gates, superposition, and entanglement to efficiently balance exploration and exploitation. Unlike gradient methods, QIEO is a global search algorithm that starts with initial random solutions, each of which mimics a quantum circuit. These circuits are being updated with the help of parametric gates based on the heuristics that are used. The circuit updation continues until it reaches convergence, i.e., reaches the optimal solution. The general workflow of QIEO is as follows:
  1. 1.
    Initialize qubit population: All qubits in equal superposition (e.g., 50-50 chance of 0 or 1).
  2. 2.
    Sample individuals from qubit states.
  3. 3.
    Evaluate the fitness of the individuals sampled.
  4. 4.
    Select the best solutions.
  5. 5.
    Update qubits using rotation rules for the best solutions.
  6. 6.
    Repeat until the stopping condition is met (max generations, fitness threshold, etc.).
The following are the inputs required for QIEO: the number of unknown qubits, the number of quantum individuals, the maximum number of generations, and the change in theta for the evolution of the gate $R_{y}$. The last parameter, that is, the change in theta, is particular to QIEO, which affects the exploration and the exploration rate. In general, QIEO provides a reliable solution that is superior in terms of quality and consistency compared to its traditional counterpart. The fact that QIEO requires a lower population to achieve better accuracy makes this algorithm suitable for solving complex optimization problems Kumar et al. (2024, 2025).
IV.2 Performance Comparison: Optimization using BQPhy’s QIEO vs. Optimization using Traditional Genetic Algorithm on a Classical Computer
To evaluate the computational efficiency of the quantum-inspired optimization framework (BQP), we compare its performance with a classical genetic algorithm approach implemented on a high-performance personal workstation. The algorithm is used to minimize the free-energy landscape and identify optimal spin configurations under uncertainty.
BQP-based optimization leverages quantum-inspired evolutionary operations to directly search high-dimensional discrete design spaces, eliminating the need for an explicit surrogate. In contrast, the classical surrogate approach benefits from reduced evaluation costs during optimization but requires a significant initial investment in training time. This section provides a comparative analysis of the computational cost associated with each approach.
All classical experiments were conducted on a high-performance cluster equipped with an AMD EPYC 7742 chip, featuring 64 cores and 128 threads with a base clock speed of 2.3 GHz and a turbo speed of 3.4 GHz. The system was outfitted with 16 GB of RAM and a 11.7 TB solid-state drive (SSD). These specifications provide a representative baseline for evaluating computational performance on modern consumer-grade hardware.
(a) BQP
(b) GA
(c) BQP
(d) GA
(e) BQP
(f) GA
Figure 3: For comparison, the optimum solutions obtained using both BQP and GA are illustrated for the $15\times 15$, $25\times 25$, and $50\times 50$ domains.
The second optimization problem given in Eq. 5 is solved with both GA and BQP for different domains and the total computational times and optimum solutions for each method are summarized in Table 1. For the traditional genetic algorithm (GA), the optimization required 164.05 seconds for a $10\times 10$ lattice. In contrast, the BQP solver completed the process in only 75.29 seconds without any loss of accuracy. Extending to larger domains, BQP consistently achieved the same accuracy in roughly half the time of GA. The optimum solutions produced by both methods are presented in Fig. 3. Since the problem explicitly accounts for uncertainty, the existence of multiple equivalent optima is expected and is indeed observed across all solutions. BQP’s direct handling of discrete binary variables may be advantageous primarily for large-scale systems with complex energy landscapes.
Table 1: Performance Comparison of BQP and GA in Minimizing $std_{f_{m}}$: BQP vs GA
Domain Method Optimum Solution Optimization Time (s)
$10\times 10$ BQP (QIEO) 0.0184008 75.29
GA 0.0184088 164.05
$15\times 15$ BQP (QIEO) 0.0096372 298.10
GA 0.0096985 565.94
$25\times 25$ BQP (QIEO) 0.0040946 1911.785
GA 0.0040969 3452.094
$50\times 50$ BQP (QIEO) 0.0012304 25600.182
GA 0.0012368 46576.803
These results highlight the promise of quantum-inspired methods such as BQP for accelerating the design of magnetic lattices. By substantially reducing computational time without compromising solution quality, BQP demonstrates scalability and efficiency that classical evolutionary approaches struggle to match. This advantage becomes increasingly significant as system size grows, suggesting that quantum-inspired optimization could play a pivotal role in tackling next-generation problems in materials design.
V Discussions and Future Work
This study demonstrates that the incorporation of uncertainty in external parameters such as temperature and magnetic field fundamentally affects the optimization landscape for spin systems. While the first problem converges almost instantly to a uniform spin configuration, the second problem requires a much larger number of iterations to approach an optimum solution. This increase in computational demand arises from the variability introduced by modeling $T$ and $h$ as Gaussian-distributed random variables and from the necessity of exploring a wide design space to capture their effects. As a result, classical algorithms such as GA become computationally expensive, especially when long-range spin interactions are also considered.
In this context, the application of BQP’s quantum-inspired evolutionary optimization offers an important alternative. By encoding discrete spin states through probabilistic amplitudes and updating them with quantum-inspired operators, BQP provides a more efficient search of the energy landscape. The comparison between GA and BQP shows that for uncertainty-aware optimization, BQP achieves comparable objective values at significantly reduced computational times. This efficiency advantage shows the potential of quantum-inspired approaches for complex Hamiltonian-based models where classical methods scale poorly. Moreover, the success of BQP in this study emphasizes that hybrid strategies, which borrow from quantum principles while running on classical high-performance computing platforms, can serve as an effective placeholder until large-scale quantum hardware becomes more accessible.
Future work will extend the present framework along several directions. First, while the current study focuses on a $10\times 10$ lattice, larger and higher-dimensional systems remain to be investigated, where the computational burden of GA may become prohibitive and the benefits of BQP more pronounced. Second, additional forms of uncertainty, such as variations in long-range weight parameters $\omega_{j}$ or lattice geometry, could be incorporated to further evaluate robustness. Third, the role of long-range interactions modeled via Gaussian distributions could be explored in more detail under uncertainty to assess their contribution to critical phenomena such as the ferromagnetic–paramagnetic phase transition. Finally, the comparative performance of GA and BQP across these more demanding scenarios will help clarify the domains in which quantum-inspired methods consistently outperform classical optimization.
Conflict of Interest Statement
On behalf of all authors, the corresponding author states that there is no conflict of interest.
Credit
The following article has been submitted to APL Quantum. After it is published, it will be found at https://publishing.aip.org/resources/librarians/products/journals/.
Acknowledgments
Virginia Tech authors would like to acknowledge the financial support from the Air Force Office of Scientific Research (AFOSR) Grants FA9550-21-1-0120 and FA9550-24-1-0154 (Program Officer: Dr. Ali Sayir) for the development of the Ising model solution and magnetic domain formulation.